# TAR3D: Creating High-Quality 3D Assets via Next-Part Prediction

### v1

@@ 4.4 Ablation Studies
We conduct ablative experiments on the chair category of the ShapeNet dataset to verify the effectiveness of our key designs in the proposed TAR3D framework, i.e., 3D VQ-VAE and 3D GPT.
@@
Evaluation on 3D VQ-VAE. We first evaluate the 3D reconstruction capability of our 3D VQ-VAE, which serves as a foundation for generating high-quality 3D assets. For reference, we also provide the performance of the VAE counterpart from which our 3D VQ-VAE is derived. As shown in Tab. 2, the proposed 3D VQ-VAE achieves excellent reconstruction performance, surpassing that of its VAE variant by about 10%. These experimental results demonstrate that our 3D VQ-VAE can encode the 3D shapes into a highly representative latent space, which enables the efficient autoregressive generation of 3D geometry parts in 3D GPT.
Table 2: Ablation studies on the capability of our VQ-VAE and its originating VAE for 3D reconstruction.
Chamfer Distance $\downarrow$ F-Score $\uparrow$
VAE 0.018 0.811
3D VQ-VAE 0.016 0.822
@@
Plane information interaction in 3D VQ-VAE. Then, we analyze the importance of plane information interaction (PII) achieved by the feature deformation and attention mechanism in our 3D VQ-VAE. As shown in Tab. 3, the variant without PII designs achieves $0.661$ f-score in 3D reconstruction. With the incorporation of our PII designs, this score significantly increased to $0.822$. These experiments indicate that our PII designs contribute to the decoding process from the latent features to 3D objects.
Table 3: Ablation studies on the plane information interaction (PII) designs in the decoder of 3D-VQVAE.
Chamfer Distance $\downarrow$ F-Score $\uparrow$
w/o PII 0.023 0.661
w/ PII 0.016 0.822
@@
TriPE positional encoding in 3D GPT. We also investigate two positional encoding strategies for the sequence modeling in 3DGPT. The first strategy is the 1D Rotary Position Embedding (RoPE) [54] matching the sequence length, and the other is the proposed TriPE. As can be observed in Fig. 8, the 1D RoPE is prone to losing important geometric details of the objects in the input images. In contrast, our well-designed TriPE can preserve as much 3D spatial information as possible, thereby generating 3D objects that are geometrically more faithful to the prompts.
Figure 8: Ablation experiments on the effectiveness of our TriPE in 3D GPT. All views are rendered at a fixed elevation of $20^{\circ}$.
5 Conclusion & Future Work
@@
In this paper, we propose a novel framework named TAR3D to generate high-quality 3D assets via the “next-token prediction” paradigm derived from multimodal large language models. To achieve this, we first develop a 3D VQ-VAE, in which the 3D shapes are encoded into the triplane latent space and quantized as discrete embeddings from a trainable codebook. Equipped with the well-designed feature deformation and attention mechanism, these quantized features are utilized to reconstruct fine-grained geometries in a neural occupancy field. Then, we adopt the codebook indices of these discrete representations to form the sequence for autoregressive modeling. Driven by the prefilling prompt embeddings, our 3D GPT, along with the 3D spatial information from the proposed TriPE, can generate high-quality 3D objects part by part. Finally, extensive experiments on ShapeNet and Objaverse, are conducted to demonstrate the superior performance of our TAR3D compared to existing 3D generation approaches.
@@
We believe it is promising to improve 3D object generation under the autoregressive setting. There are a few foreseeable directions to explore: 1) Scaling law. We will collect more 3D data with prompts from other datasets, e.g., Objaverse-XL [16], 3D-FUTURE [20], and scale the GPT model for a stronger generation capability. 2) Sequence formulation. We will explore the sequence formulation manner of triplane indices for more efficient generative modeling. 3) Tokenization. We will establish a unified codebook for 3D representations and prompt embeddings.

### v2

@@ 4.3 Quantitative Evaluation
Table 2: Ablation studies on the reconstruction capability of the 3D VQ-VAE.
3D VAE 3D VQ-VAE
Chamfer Distance $\downarrow$ 0.018 0.016
F-Score $\uparrow$ 0.811 0.822
Table 3: Ablation studies on the PII design of the 3D-VQVAE.
w/o PII w/ PII
Chamfer Distance $\downarrow$ 0.023 0.016
F-Score $\uparrow$ 0.661 0.822
Table 4: Ablation studies on triplane size (i.e., sequence length) of the 3D GPT.
Triplane Size $3\!\times\!16\!\times\!16$ $3\!\times\!32\!\times\!32$ $3\!\times\!48\!\times\!48$
Chamfer Distance $\downarrow$ 0.157 0.066 0.062
Inference Time $\downarrow$ 17.7 s 67.6 s 143.9 s
@@ 4.4 Ablation Studies
Evaluation on 3D VQ-VAE. We evaluate the 3D reconstruction capability of our 3D VQ-VAE, which serves as a foundation for generating high-quality 3D assets. For reference, we provide the performance of the VAE counterpart from which our 3D VQ-VAE is derived. To obtain a well-performing tokenizer for autoregressive generation, we adjust the training strategies until the performance of our 3D VQ-VAE is on par with or better than that of the VAE counterpart. The experiment results shown in Tab. 4 demonstrate the reconstruction effectiveness of our 3D VQ-VAE.
@@
Plane information interaction in 3D VQ-VAE. We analyze the importance of plane information interaction (PII) achieved by the feature deformation and attention mechanism in our 3D VQ-VAE. As shown in Tab. 4, the variant without PII design achieves $0.661$ f-score in 3D reconstruction. With the incorporation of our PII designs, this score significantly increased to $0.822$. These experiments indicate that our PII designs contribute to the decoding process from the latent features to 3D objects.
@@
TriPE positional encoding in 3D GPT. We investigate two positional encoding strategies for the sequence modeling in 3DGPT. The first strategy is the 1D Rotary Position Embedding (RoPE) [56] matching the sequence length, and the other is the proposed TriPE. As can be observed in Fig. 6, the 1D RoPE is prone to lose important geometric details of the objects in the input images. In contrast, our well-designed TriPE can preserve as much 3D spatial information as possible, thereby generating 3D objects that are geometrically more faithful to the prompts.
Figure 6: Ablation experiments on the effectiveness of our TriPE.
Figure 7: Visual comparisons of our TAR3D variants with different triplane sizes or sequence lengths.
Triplane Size. We ablate the triplane size with 3$\times$16$\times$16, 3$\times$32$\times$32, and 3$\times$48$\times$48 to explore the impact of sequence length on the performance and efficiency. Note that the 3$\times$ 32$\times$32 widely adopted in 3D reconstruction is our default setting. As shown in Tab. 4, it is the best setting for efficiency and performance trade-offs, which is also demonstrated by the visual comparisons in Fig. 7
5 Conclusions & Future Work
@@
In this paper, we propose a novel framework named TAR3D to generate high-quality 3D assets via the “next-token prediction” paradigm derived from multimodal large language models. To achieve this, we first develop a 3D VQ-VAE, in which the 3D shapes are encoded into the triplane latent space and quantized as discrete embeddings from a trainable codebook. Equipped with the well-designed feature deformation and attention mechanism, these quantized features are utilized to reconstruct fine-grained geometries in a neural occupancy field. Then, we adopt the codebook indices of these discrete representations to form the sequence for autoregressive modeling. Driven by the prefilling prompt embeddings, our 3D GPT, along with the 3D spatial information from the proposed TriPE, can generate high-quality 3D objects part by part. Finally, extensive experiments on several benchmark and out-of-domain 3D datasets, are conducted to demonstrate the superior performance of our TAR3D compared to existing approaches.
@@
We believe it is promising to improve 3D object generation under the autoregressive setting. There are a few foreseeable directions to explore: 1) Scaling law. We will collect more 3D data with prompts from other datasets, e.g., Objaverse-XL [17], 3D-FUTURE [22], and scale the GPT model for a stronger generation capability. 2) Sequence formulation. We will explore the sequence formulation manner of triplane indices for more efficient generative modeling. 3) Tokenization. We plan to establish a unified codebook for 3D representations and prompt embeddings.
[…]
Supplementary Material
Appendix A More Visualization Results
Figure 8: Given a wide range of input images from the benchmark datasets, i.e., ShapeNet and Objaverse, our TAR3D model is capable of inferring the full picture of the primary objects and producing high-quality 3D meshes that adhere to them.
Figure 9: Given a wide range of input images from the out-of-domain dataset, i.e., GSO, our TAR3D model is capable of inferring the full picture of the primary objects and producing high-quality 3D meshes that adhere to them.
Apart from the quantitative and qualitative comparisons of the proposed TAR3D with recent representative conditional 3D object generation approaches in the main text, in this section, we also provide more synthesized 3D meshes of TAR3D under image-to-3d and text-to-3d settings.
Fig. 8 and Fig. 9 show the generated meshes when feeding our TAR3D with image prompts from the benchmark datasets (Objaverse and ShapeNet) and the out-of-domain dataset (GSO), respectively. Our TAR3D can infer the full picture of an object from specific views, resulting in 3D assets that adhere to the given images.
Fig. 10 illustrates the 3D objects created by our TAR3D according to various textual descriptions. Note that these text prompts come from the ULIP dataset [71], in which the texts are utilized to describe the rendered images while containing some noise for 3D objects, e.g., “a grey background” in the sentence describing the camera model. Nevertheless, our TAR3D can still generate faithful 3D assets for the primary objects described in these sentences.
Appendix B More Discussions about Future Work
Although this work establishes an autoregressive 3D generation pipeline based on triplane representations from scratch, it still has certain limitations. We summarize them as follows and leave them for future research.
  • Data Cleaning and Expansion. The prompt system of our TAR3D is built on the ULIP dataset [71]. However, the quality of the rendered images is unsatisfactory, and there is a certain level of noise in the textual descriptions. We will re-render the collected 3D objects and adopt the Cap3D strategy [44] to caption them. In addition, we will collect more data with prompts from other 3D datasets, like Objaverse-XL [17], Omniobject3d [69], ABO [15], etc.
  • Model Scaling Up. The 3D GPT architecture is almost the same as Llama. Once the aforementioned 3D dataset is assembled, it is promising to scale the GPT model via optimization techniques and training recipes in the LLM community for a stronger 3D generation capability.
  • Sequence formulation. The sequence length in this work is three times that of image autoregressive generation, as the triplane representation consists of three feature maps. Although shorter than methods based on mesh face quantization, it still has room for improvement. In future work, we will explore more concise sequence formulations of triplane indices for efficient autoregressive modeling.
  • Tokenization. In this paper, we employ the encoded image or text features as prefilling token embeddings to integrate such conditions into autoregressive models. Considering the powerful multimodal unification capability of the GPT model, we will establish a unified codebook for 3D representations and prompt embeddings.